### v1

@@ Abstract
Dialectal Arabic (DA) poses a persistent challenge for natural language processing (NLP), as most everyday communication in the Arab world occurs in dialects that diverge significantly from Modern Standard Arabic (MSA). This linguistic divide limits access to digital services and educational resources and impedes progress in Arabic machine translation. This paper presents two core contributions to advancing DA–MSA translation for the Levantine, Egyptian, and Gulf dialects, particularly in low-resource and computationally constrained settings: a comprehensive evaluation of training-free prompting techniques, and the development of a resource-efficient fine-tuning pipeline. Our evaluation of prompting strategies across six large language models (LLMs) found that few-shot prompting consistently outperformed zero-shot, chain-of-thought, and our proposed Ara-TEaR method. GPT-4o achieved the highest performance across all prompting settings. For fine-tuning, a quantized Gemma2-9B model achieved a CHrF++ score of 49.88, outperforming zero-shot GPT-4o (44.58). Joint multi-dialect trained models outperformed single-dialect counterparts by over 10% CHrF++, and 4-bit quantization reduced memory usage by 60% with less than 1% performance loss. The results and insights of our experiments offer a practical blueprint for improving dialectal inclusion in Arabic NLP, showing that high-quality DA–MSA machine translation is achievable even with limited resources and paving the way for more inclusive language technologies.
@@
Machine Translation Dialectal Arabic Modern Standard Arabic Translation Evaluation Large Language Model Fine-Tuning Prompting Technique
@@ 1 Introduction
Arabic is the fifth most spoken language globally, with over 420 million speakers across more than 20 countries Eberhard et al (2023). This vast linguistic landscape is marked by a significant linguistic divide: Modern Standard Arabic (MSA) serves as the formal register, while numerous dialectal varieties, such as Egyptian, Levantine, and Gulf, dominate everyday communication. These dialects diverge considerably from MSA in vocabulary, style, and syntax, posing substantial barriers to language technologies. Existing natural language processing (NLP) systems, while effective for MSA, often fail when faced with dialectal Arabic (DA) content. For example, social media monitoring tools, built primarily for MSA, struggle to interpret sentiment and meaning in dialectal content, which constitutes most of the online user-generated text of Arabic Abdul-Mageed et al (2020).
The recent emergence of Large Language Models (LLMs) offers promising new directions for Arabic Machine Translation (MT). However, research addressing specifically the application of LLMs to the translation of DA-MSA remains notably limited Farhan et al (2023).
Although some efforts have focused on DA-MSA MT methods, there remains a critical need for two specific areas of investigation: (1) a comprehensive and systematic evaluation of diverse prompting techniques across various LLMs for DA-MSA translation, and (2) a detailed study featuring a systematic experimentation process to identify performance-critical variables affecting translation quality when fine-tuning LLMs for low-resource MT.
This paper addresses these gaps by extensively exploring and evaluating two distinct approaches for DA-MSA neural machine translation within low-resource and computationally constrained environments. The first approach investigates training-free prompting strategies, systematically assessing how different prompt designs and contextual cues impact translation quality across various dialects and LLMs without the need for model fine-tuning. The second approach focuses on developing resource-efficient fine-tuning methodologies that achieve high-quality DA-MSA translation while significantly reducing computational requirements. This involves exploring techniques such as Low-Rank Adaptation (LoRA), quantization, data augmentation, and multi-dialect joint training, with systematic comparisons presented to identify the trade-offs between translation quality and efficiency.
@@ 2 Background and Related Work
2.1 Traditional Approaches to DA-MSA translation
@@
In addition to rule-based transformation, lexical mapping and transliteration strategies were explored. Dictionary-based lookups and transliteration tools sought to align phonetically similar tokens, but were hindered by high orthographic variability and lack of standardized spelling in dialects (Zaghouani et al, 2018).
Phrase-based statistical machine translation (SMT), as implemented in systems like Moses (Koehn et al, 2007), represented a major step forward. SMT systems used probabilistic phrase alignments and required bilingual corpora to learn translation patterns. However, the scarcity of high-quality DA–MSA parallel corpora limited SMT performance (Abu Farha and Magdy, 2020). Researchers attempted to compensate by generating synthetic parallel data or leveraging corpora such as MADAR and BTEC, but the lack of broad dialectal coverage remained a bottleneck.
@@
2.2 Neural Machine Translation
@@
Neural Machine Translation (NMT) significantly advanced the capabilities of DA-MSA translation systems, addressing challenges that traditional MT systems failed to handle such as the morphological richness, orthographic ambiguity, and dialectal variation of Arabic dialects Baniata et al (2021). NMT revolutionized the field by leveraging deep neural networks capable of modeling non-linear and context-dependent relationships between dialectal inputs and MSA outputs, enabling more fluent and accurate outputs across dialects  Bahdanau et al (2015); Vaswani et al (2017)
A neural network (NN) is a multi‑layer architecture of interconnected neurons trained to learn representations from data, forming the foundation of modern language models Vaswani et al (2017). NMT typically uses encoder–decoder architectures, where an encoder processes source text and a decoder generates the target translation, or decoder‑only models, which autoregressively generate each output token conditioned on previously generated tokens. While encoder–decoder models initially dominated MT, decoder‑only models have gained popularity due to their scalability and strong generative performance at scale Brown et al (2020).
@@
Compared to traditional methods, NMT offers better generalization to unseen inputs, learns complex dialect-to-MSA mappings without hand-crafted rules, and supports joint multilingual training across dialects Bahdanau et al (2015); Johnson et al (2017); Baniata et al (2021).
2.3 The emergence of LLMs
@@
The emergence of large language models (LLMs) has transformed machine translation for low-resource languages, including the understudied task of DA–MSA translation. LLMs, such as GPT and LLaMA, are transformer-based architectures trained on vast multilingual corpora to model complex linguistic patterns Vaswani et al (2017); Brown et al (2020). These models learn to predict the next token in a sequence based on the context, allowing them to generate coherent and contextually appropriate outputs across diverse linguistic inputs. Their ability to capture long-range dependencies and semantic nuances enables them to perform a wide range of NLP tasks, including translation, summarization, and question answering, with minimal task-specific training. Unlike traditional MT systems, LLMs can generalize across domains and dialects, enabling utilisation through both model prompting and fine-tuning. Model prompting involves task-specific instructions without modifying the model’s internal parameters, while fine-tuning adapts the model by updating its weights on a labeled dataset. Prompting excels in scenarios where high-quality resources are absent, whereas fine-tuning is preferred for robust, domain-adapted performance using relatively small datasets Chung et al (2022); Farhan et al (2023).
@@
Model prompting leverages in-context learning (ICL), where models generate task outputs based solely on examples and instructions provided at inference time, without parameter updates. This training-free paradigm relies on prompt engineering to achieve high-quality translations by framing tasks clearly and contextually Brown et al (2020). Effective prompts typically define the translation task, specify the source dialect, and may include exemplar translations (few-shot prompting) or structured reasoning instructions (Chain-of-Thought prompting) to guide the model Wei et al (2022).
@@
Prompting has become a central theme in LLM research due to its scalability and simplicity Liu et al (2023). For low-resource MT tasks like DA–MSA translation, prompting is promising due to its minimal data requirements and capacity to generalize across underrepresented dialects Chung et al (2022). Recent work has applied prompting to DA–MSA tasks, including TARJAMAT Abdul-Mageed et al (2023), which showed LLMs outperforming many commercial systems in Arabic dialect-to-English translation. In NADI 2024 Abdul-Mageed et al (2024), Arabic Train achieved top performance using Jais with a combined zero/few-shot setup Harrat et al (2024), while the winner in the OSACT 2024 shared task  Habash et al (2024) used ChatGPT with 3-shot prompting and self-correction to surpass their fine-tuned baseline Atwany et al (2024). Nevertheless, comparative and comprehensive evaluations of various prompting strategies across different LLMs, model sizes, and dialects remain limited. Additionally, key factors such as the impact of dialect identification, contextual cues, and model scaling on translation quality remains underexplored. All of which are areas that we aim to investigate in our work.
@@
While prompt engineering guides model outputs, fine-tuning directly adapts a pre-trained model by updating its parameters on a labelled and task-specific dataset, thereby improving its performance through specialization Howard and Ruder (2018). This process typically involves initializing an LLM with pretrained weights and continuing training on a smaller supervised corpus, allowing the model to internalize task-specific linguistic patterns and vocabulary Raffel et al (2020). Fine-tuning is particularly suitable for DA–MSA translation, as it enables dialect adaptation despite limited labelled data Chung et al (2022); Farhan et al (2023). Compared to prompting, which relies on inference-time instructions, fine-tuning often yields more robust and consistent results when even modest amounts of annotated data are available Liu et al (2023).
@@
Early work by Khalifa et al. Khalifa et al (2021) demonstrated the benefits of fine-tuning for dialectal adaptation in foundational NLP tasks. More recently, Sirius_Translators fine-tuned the encoder–decoder model AraT5 across five Arabic dialects Alahmari et al (2024), while CUFE adapted the decoder-only LLaMA-3 model with strong results Ibrahim (2024). Lahjawi-D2MSA proposed a multilingual approach using Kuwain-1.5B to translate 15 dialects with high human-rated fluency Hossain et al (2025). Despite these advancements, there is still a lack of systematic studies investigating the performance-critical variables that impact fine-tuning for DA–MSA MT.
@@
To address this, our work conducts a series of controlled ablation experiments isolating the effects of architecture, training strategy, and hyperparameters. Additionally, we evaluate the effectiveness of paraphrasing-based data augmentation, joint-dialect training, and active learning. We further analyze the trade-offs of 4-bit quantization on translation quality and efficiency. Finally, we compare fine-tuned and zero-shot prompted models across multiple architectures and sizes, offering a comprehensive view of the performance and scalability trade-offs in DA–MSA LLM-based translation.
3 Methods and Experimental Setup
@@
The rationale for employing a dual approach stems from both practical constraints and linguistic complexities, as discussed in the background section. Prompt-based methods yield competitive performance with minimal resource demands, making them suitable for settings where computational capabilities or annotated data are limited. However, such methods often face limitations in addressing the morphological complexity and dialectal variability inherent in Arabic, particularly when faced with highly informal, idiomatic, or structurally complex input not encountered during pre-training. In contrast, fine-tuning enables models to learn directly from DA–MSA parallel data, thereby producing more robust and dialect-aware translations. However, this advantage comes with an increase in training overhead and computational costs. Nevertheless, it is important to note that the inference costs associated with larger models, such as those used in prompt-based approaches, can lead to higher CO2 emissions per user request compared to smaller, fine-tuned models Strubell et al (2019). Given the methodological divergence of the two approaches, each is addressed through a tailored experimental setup, as outlined in  3.3 and  3.4
@@ 3.1 Dataset Curation
This study used a curated set of high-quality dialect-to-MSA parallel corpora to support both prompting and fine-tuning experiments. The selected data sets span both formal and informal domains, introducing variation in linguistic style and syntactic complexity. This diversity aimed to enhance the robustness and generalizability of the model across a broad spectrum of translation scenarios. The corpora encompassed three major Arabic dialects: Levantine (LEV), Gulf (GLF), and Egyptian (EGY).
The primary resource utilized was the Multi-Arabic Dialect Applications and Resources (MADAR) corpus Bouamor et al (2018), comprising approximately 43,000 parallel sentences in its training split. This travel-domain dataset, translated by professional translators, is widely recognized for its linguistic precision and consistency. MADAR has been extensively used in shared tasks such as OSACT  Habash et al (2024) and NADI Abdul-Mageed et al (2024). To complement the formal domain focus of MADAR, the Dial2MSA-Verified corpus Khered et al (2025) (19,775 training pairs) was also incorporated. This dataset features social media content characterized by colloquial and idiomatic expressions, and was selected due to its native-speaker validation, supporting broader domain adaptation.
To further improve the models’ understanding of Modern Standard Arabic, 5,000 MSA–MSA sentence pairs from the UFAL Parallel Corpus of North Levantine Arabic Krubiński et al (2023) were included. These same-source-target examples served as a form of regularization, enabling the model to learn when no transformation is necessary, particularly relevant given the structural affinities between Arabic dialects and MSA. In addition, a synthetic dataset of 43,000 samples was constructed by prompting GPT-4o to paraphrase the MSA references from the MADAR training data. This augmentation aims to increase lexical and stylistic variation on the MSA side while preserving the original dialectal input. Paraphrasing was preferred over back-translation due to its capacity to enrich lexical diversity without altering the dialectal text, an important consideration given the morphological complexity of Arabic. Furthermore, manual testing revealed that existing language models underperformed in translating from MSA to Dialectal Arabic, making back-translation suboptimal. The total cost of generating synthetic data via the OpenAI API OpenAI (2024) was £36.
@@
  • D1 included only the MADAR training split.
  • D2 extended D1 with Dial2MSA-Verified training data.
  • D3 further incorporated the UFAL MSA–MSA pairs and the GPT-4o-generated synthetic data.
@@
Dataset LEV GLF EGY Total
D1 17k 13.8k 12.2k 43k
D2 21.1k 20.4k 21.3k 62.8k
D3 38.1k 34.2k 33.5k 110.8k
Table 1: Breakdown of training datasets D1–D3 by dialect and size.
@@
  • T1 comprised 600 examples (200 per dialect), sampled from the test split of MADAR.
  • T2 consisted of 1,200 examples (400 per dialect), drawn from the test split of both MADAR (70%) and Dial2MSA-Verified (30%) to ensure domain and linguistic variability.
Test Split LEV GLF EGY Total
T1 200 200 200 600
T2 400 400 400 1200
Table 2: Breakdown of test datasets T1 & T2 by dialect and size.

### v2

@@ Abstract
Dialectal Arabic (DA) poses a persistent challenge for natural language processing (NLP), as most everyday communication in the Arab world occurs in dialects that diverge significantly from Modern Standard Arabic (MSA). This linguistic divide impedes progress in Arabic machine translation. This paper presents two core contributions to advancing DA–MSA translation for the Levantine, Egyptian, and Gulf dialects, particularly in low-resource and computationally constrained settings: (i) a comprehensive evaluation of training-free prompting techniques, and (ii) the development of a resource-efficient fine-tuning pipeline. Our evaluation of prompting strategies across six large language models (LLMs) found that few-shot prompting consistently outperformed zero-shot, chain-of-thought, and our proposed Ara-TEaR method. Ara-TEaR is designed as a three-stage self-refinement prompting process, targeting frequent meaning-transfer and adaptation errors in DA–MSA translation. In this evaluation, GPT-4o achieved the highest performance across all prompting settings. For fine-tuning LLMs, a quantized Gemma2-9B model achieved a chrF++ score of 49.88, outperforming zero-shot GPT-4o (44.58). Joint multi-dialect trained models outperformed single-dialect counterparts by over 10% chrF++, and 4-bit quantization reduced memory usage by 60% with less than 1% performance loss. The results and insights of our experiments offer a practical blueprint for improving dialectal inclusion in Arabic NLP, showing that high-quality DA–MSA machine translation is achievable even with limited resources and paving the way for more inclusive language technologies.
@@
Machine Translation Dialectal Arabic Modern Standard Arabic Translation Evaluation Large Language Model Fine-Tuning
@@ 1 Introduction
Arabic is the fifth most spoken language globally, with over 334 million speakers (Eberhard et al, 2023). This vast linguistic landscape is marked by a significant linguistic divide: Modern Standard Arabic (MSA) serves as the formal register, while numerous dialectal varieties, such as Egyptian, Levantine, and Gulf, dominate everyday communication. These dialects diverge considerably from MSA in vocabulary, style, and syntax, posing substantial barriers to language technologies. Existing natural language processing (NLP) systems, while effective for MSA, often fail when faced with dialectal Arabic (DA) content (Hadj Mohamed et al, 2023). For example, social media monitoring tools, built primarily for MSA, struggle to interpret sentiment and meaning in dialectal content, which is widely used in everyday communication and increasingly prevalent in social media text (Alhafni et al, 2024).
Despite the rapid progress of Large Language Models (LLMs), research specifically applying them to effectively translate dialectal Arabic into Modern Standard Arabic (DA–MSA) remains notably limited. Benchmarking efforts like AraDiCE (Mousi et al, 2025) highlight persistent gaps in dialect translation and comprehension. Although some efforts have focused on DA-MSA MT methods, there remains a critical need for two specific areas of investigation: (1) a comprehensive and systematic evaluation of diverse prompting techniques across various LLMs for DA-MSA translation, and (2) a detailed study featuring a systematic experimentation process to identify performance-critical variables affecting translation quality when fine-tuning LLMs for low-resource DA-MSA MT.
This paper addresses these gaps by extensively exploring and evaluating two distinct approaches for DA-MSA neural machine translation (NMT) within low-resource and computationally constrained environments. The first approach investigates training-free prompting strategies, systematically assessing how different prompt designs and contextual cues impact translation quality across various dialects and LLMs without the need for model fine-tuning. The second approach focuses on developing resource-efficient fine-tuning methodologies that achieve high-quality DA-MSA translation while significantly reducing computational requirements. This involves exploring techniques such as Low-Rank Adaptation (LoRA), quantization, data augmentation, and multi-dialect joint training, with systematic comparisons presented to identify the trade-offs between translation quality and efficiency.
To support future research and development, our data and code are openly shared at Github  [^2] https://github.com/er-abd/Advancing-DA2MSA-MT. Our fine-tuned models can be downloaded from Huggingface  [^3] https://huggingface.co/collections/er-abd/ (advancing-da2msa-mt).
@@ 2 Background and Related Work
2.1 Traditional Approaches for DA-MSA translation
@@
In addition to rule-based transformation, lexical mapping and transliteration strategies were explored. Dictionary-based lookup and transliteration tools sought to align phonetically similar tokens, but were hindered by high orthographic variability and lack of standardized spelling in dialects (Habash et al, 2012).
Phrase-based statistical machine translation (SMT), as implemented in systems like Moses (Koehn et al, 2007), represented a major step forward. SMT systems used probabilistic phrase alignments and required bilingual corpora to learn translation patterns. However, the scarcity of high-quality DA–MSA parallel corpora has limited SMT performance (Klementiev et al, 2012). Researchers attempted to compensate by generating synthetic parallel data or leveraging existing corpora such as MADAR (Bouamor et al, 2018), but the lack of broad dialectal coverage remained a bottleneck.
@@
2.2 Neural Machine Translation for DA-MSA translation
@@
Neural Machine Translation (NMT) significantly advanced the capabilities of DA-MSA translation systems, addressing challenges that traditional MT systems failed to handle such as the morphological richness, orthographic ambiguity, and dialectal variation of Arabic dialects (Baniata et al, 2021). NMT revolutionized the field by leveraging deep neural networks capable of modeling non-linear and context-dependent relationships between dialectal inputs and MSA outputs, enabling more fluent and accurate translation across dialects  (Vaswani et al, 2017)
@@
A neural network (NN) is a multi‑layer architecture of interconnected neurons trained to learn representations from data, forming the foundation of modern language models (LeCun et al, 2015). NMT typically uses encoder–decoder architectures, where an encoder processes source text and a decoder generates the target translation. Compared to traditional methods, NMT offers better generalization to unseen inputs, learns complex dialect-to-MSA mappings without hand-crafted rules, and supports joint multilingual training across dialects (Johnson et al, 2017; Baniata et al, 2021).
2.3 Prompting and Fine-Tuning LLMs for DA–MSA Translation
@@
The decoder-only architecture has proven highly effective for MT, as it autoregressively generates each output token conditioned on prior tokens. While encoder–decoder models initially dominated, decoder-only models gained popularity for their scalability and strong generative performance at scale (Caillaut et al, 2024). Most large language models (LLMs), including GPT and LLaMA, adopt this decoder-only Transformer architecture, trained on vast multilingual corpora to capture complex linguistic patterns (Naik et al, 2024). The emergence of LLMs has transformed MT for low-resource languages, including the understudied task of DA–MSA translation. These models learn to predict the next token in a sequence based on the context, allowing them to generate coherent and contextually appropriate outputs across diverse linguistic inputs. Their ability to capture long-range dependencies and semantic nuances enables them to perform a wide range of NLP tasks, including translation, summarization, and question answering, with minimal task-specific training. Unlike traditional MT systems, LLMs can generalize across domains and dialects, enabling utilization through both model prompting and fine-tuning. Model prompting involves task-specific instructions without modifying the model’s internal parameters, while fine-tuning adapts the model by updating its weights on a labeled dataset. Prompting excels in scenarios where high-quality resources are absent, whereas fine-tuning is preferred for robust, domain-adapted performance using relatively small datasets (Chen et al, 2023).
@@
LLM prompting leverages in-context learning (ICL), where models generate task outputs based solely on examples and instructions provided at inference time, without parameter updates. This training-free paradigm relies on prompt engineering to achieve high-quality translations by framing tasks clearly and contextually (Liu et al, 2023). Effective prompts typically define the translation task, specify the source dialect, and may include exemplar translations (few-shot prompting) or structured reasoning instructions (Chain-of-Thought prompting) to guide the model (Wei et al, 2022).
@@
Prompting has become a central theme in LLM research due to its scalability and simplicity (Liu et al, 2023). For low-resource MT tasks like DA–MSA translation, prompting is promising due to its minimal data requirements and capacity to generalize across underrepresented dialects (Chung et al, 2024). Recent work has applied prompting to DA–MSA translation, including TARJAMAT (Kadaoui et al, 2023), which showed LLMs outperforming many commercial systems in Arabic dialect-to-English translation. In NADI 2024 (Abdul-Mageed et al, 2024), Arabic Train achieved top performance using Jais with a combined zero/few-shot setup (Demidova et al, 2024), while the winner in the OSACT 2024 shared task  (Al-Khalifa et al, 2024) used ChatGPT with 3-shot prompting and self-correction to surpass their fine-tuned baseline (Atwany et al, 2024). Nevertheless, comparative and comprehensive evaluations of various prompting strategies across different LLMs, model sizes, and dialects remain limited. Additionally, key factors such as the impact of dialect identification, contextual cues, and model scaling on translation quality remains underexplored, all of which are areas that we aim to investigate in our work.
@@
While prompt engineering guides model outputs, fine-tuning directly adapts a pre-trained model by updating its parameters on a labeled and task-specific dataset, thereby improving its performance through specialization (Howard and Ruder, 2018). This process typically involves initializing an LLM with pretrained weights and continuing training on a smaller supervised corpus, allowing the model to internalize task-specific linguistic patterns and vocabulary (Raffel et al, 2020), e.g. in MT (Yin Hong et al, 2024). Fine-tuning is particularly suitable in low-resource settings like DA–MSA translation, as it enables models to adapt to specific domains with limited labeled data. Compared to prompting, which relies on inference-time instructions, fine-tuning often yields more robust and consistent results when even modest amounts of annotated data are available (Liu et al, 2023; Han et al, 2024).
@@
In Arabic domain, an early work by Khalifa et al (2021) demonstrated the benefits of fine-tuning for dialectal adaptation in foundational NLP tasks. More recently, Sirius_Translators fine-tuned the encoder–decoder model AraT5 across five Arabic dialects (Alahmari, 2024), while CUFE adapted the decoder-only LLaMA-3 model with strong results (Ibrahim, 2024). Lahjawi-D2MSA, a fine-tuned version of the Kuwain-1.5B small language model, was proposed for translating 15 Arabic dialects to Modern Standard Arabic and achieved a high human-rated fluency score of 78% in fluency evaluations (Hamed et al, 2025). Despite these advancements, there is still a lack of systematic studies investigating the performance-critical variables that impact fine-tuning for DA–MSA MT.
@@
To address this, our work conducts a series of controlled ablation experiments isolating the effects of architecture, training strategy, and hyper-parameters. Additionally, we evaluate the effectiveness of paraphrasing-based data augmentation, joint-dialect training, and active learning. We further analyze the trade-offs of 4-bit quantization on translation quality and efficiency. Finally, we compare fine-tuned and zero-shot prompted models across multiple architectures and sizes, offering a comprehensive view of the performance and scalability trade-offs in DA–MSA LLM-based translation.
3 Methodology
@@
The rationale for employing a dual approach stems from both practical constraints and linguistic complexities, as discussed in Section  2. Prompt-based methods yield competitive performance with minimal resource demands, making them suitable for settings where computational capabilities or annotated data are limited. However, such methods often face limitations in addressing the morphological complexity and dialectal variability inherent in Arabic, particularly when faced with highly informal, idiomatic, or structurally complex input not encountered during pre-training  (Kadaoui et al, 2023). In contrast, fine-tuning enables models to learn directly from DA–MSA parallel data, thereby producing more robust and dialect-aware translations. However, this advantage comes with an increase in training overhead and computational costs. Nevertheless, it is important to note that the inference costs associated with larger, general-purpose models used in prompting approaches can lead to significantly higher carbon emissions per user request than smaller, fine-tuned models, even with similar parameter counts (Luccioni et al, 2024). Given the methodological divergence of the two approaches, each is addressed through a tailored experimental setup, as outlined in  3.3 and  3.4
@@ 3.1 Dataset Curation
This study used a curated set of high-quality DA-to-MSA parallel corpora to support both prompting and fine-tuning experiments. The selected datasets span both formal and informal domains, introducing variation in linguistic style and syntactic complexity. This diversity aimed to enhance the robustness and generalizability of the model across a broad spectrum of translation scenarios. Our curated corpora encompassed three major Arabic dialects: Levantine (LEV), Gulf (GLF), and Egyptian (EGY).
The primary resource utilized was the Multi-Arabic Dialect Applications and Resources (MADAR) corpus (Bouamor et al, 2018), comprising approximately 43,000 parallel sentences in its training split. This travel-domain dataset, translated by professional translators, is widely recognized for its linguistic precision and consistency. MADAR has been extensively used in shared tasks such as OSACT  (Al-Khalifa et al, 2024) and NADI (Abdul-Mageed et al, 2024). To complement the formal domain focus of MADAR, the Dial2MSA-Verified corpus (Khered et al, 2025) (19,775 training pairs) was also incorporated. This dataset features social media content characterized by colloquial and idiomatic expressions, and was selected due to its native-speaker validation.
To further improve the models’ understanding of MSA, 5,000 MSA–MSA sentence pairs from the UFAL Parallel Corpus of North Levantine Arabic (Krubiński et al, 2023) were included. These same-source-target examples served as a form of regularization, enabling the model to learn when no transformation is necessary, particularly relevant given the structural affinities between Arabic dialects and MSA. In addition, a synthetic dataset of 43,000 samples was constructed by prompting GPT-4o to paraphrase the MSA references from the MADAR training data. This augmentation aims to increase lexical and stylistic variation on the MSA side while preserving the original dialectal input. Paraphrasing was preferred over back-translation due to its capacity to enrich lexical diversity without altering the dialectal text, an important consideration given the morphological complexity of Arabic. Furthermore, manual testing revealed that existing language models underperformed in translating from MSA to Dialectal Arabic, making back-translation suboptimal.
@@
  • Minimal-Gold included only the MADAR training split.
  • Extended-Gold extended the Minimal-Gold dataset with Dial2MSA-Verified training data.
  • Extended-Gold-Silver extended the Extended-Gold (62,775 examples) with 43,000 GPT-4o–paraphrased MSA references and 5,000 UFAL MSA–MSA pairs, yielding a total of 110,775 training examples.
@@
Training Set LEV GLF EGY Total
Minimal-Gold 17k 13.8k 12.2k 43k
Extended-Gold 21.1k 20.4k 21.3k 62.8k
Extended-Gold-Silver 38.1k 34.2k 33.5k 110.8k
Table 1: Breakdown of training datasets Minimal-Gold , Extended-Gold , and Extended-Gold-Silver by dialect and size
@@
  • Small-Test comprised 600 examples (200 per dialect), randomly sampled from the test split of MADAR.
  • Large-Test comprised of 1,200 examples (400 per dialect), randomly sampled from the test split of both MADAR (70%) and Dial2MSA-Verified (30%) to ensure domain and linguistic variability.
Test Set LEV GLF EGY Total
Small-Test 200 200 200 600
Large-Test 400 400 400 1200
Table 2: Breakdown of test datasets Small-Test & Large-Test by dialect and size
The two test sets might contain potential overlapping examples since both included randomly sampled examples from the test split of MADAR.